OpenStudy (anonymous):

David planted 18 plants. Five sixth of the plants are tomato plants. How many of the plants are tomato plants? Is it 15?

OpenStudy (anonymous):

\[\frac{ 18 }{ 1 } * \frac{ 5 }{ 6 } = \frac{ 90 }{ 6 } = 15\]

OpenStudy (anonymous):

Yes it is 15 :)